Operations Report – September 6th

MDRS Complete Operations Report 06-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action is taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 79.7
Minimum 62
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – not checked gallons

Water (loft tank): 50 gallons
Water Meter: n/a units
Water (static tank) – 150 gallons
Static to Loft Pump used – yes
Water in GreenHab – 175 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: yes

Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge:
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): All 350’s used
Reason for use: getting rocks for the tourist kiosk, geology
Oil Added? No
ATV Fuel Used: Not checked Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, travel back and forth from town storage units

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are charging.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: We built the pathway for the new tourist kiosk today. We also constructed part of the new tunnel coverings.

We did another purge of the storage units, this time going through and removing materials left over from the construction of the Hab! We also decided to remove things we had been storing for past crews, researchers, and expeditions. These are things that had been unclaimed for years.

I’m happy to report that much of this is finding new homes in Hanksville. We are still far from finished with the work there, but it’s a good start. The mud is dried except for where it has been trapped under or inside things but it smells really bad.

The goal is to remove everything from one unit, get what remakes all on shelves in a single unit, and then cleaned the floors of both. There will be room for storage of some things with the URC materials but I would like to get all things of use back at MDRS. For example, all of our power tools were in there but we could not access them.

Most of the damaged Ikea material was removed from packaging and cleaned and dried. We will be able to use most of it still.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Prep has begun for the new wall in the intern trailer bedroom. We have begun troubleshooting the intern trailer plumbing.

Summary of GreenHab operations: Water prep of the soil is underway. It’s still very hot to start seeding but I think it can be done. A few large herbs have been planted.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – September 5th

MDRS Complete Operations Report 05-SEPT-2021
Name of person filing report: Shannon Rupert
Reason for Report: Beginning of field season/work party daily update
Non-nominal systems: Old leak in the kitchen sink in Hab has resurfaced in the ground below the kitchen
Action taken for non-nominal systems: Nothing to report
Generator: Nothing to report
ScienceDome Dual Split: Operating nominally and I love having AC to escape to when it’s so hot in the afternoon
Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 76.5
Minimum 49
Maximum 100
Note on solar: work party crews are not power conserving, and the heat is adding to our power consumption
Diesel Reading – 0 %.
Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%
Ethanol Free Gasoline – not checked gallons
Water (loft tank): 50 gallons
Water Meter: n/a unit
Water (static tank) – 100 gallons
Static to Loft Pump used – yes
Water in GreenHab – 200 gallons, filled
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 200 gallons
Hab toilet tank emptied: no
Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge:
Currently charging: yes
Notes on rovers: Nothing to report
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). One of the 350’s is being used as a signReason for use: n/aOil Added? NoATV Fuel Used: Not checked Gals# Hours the ATVs were Used today: 0 Notes on ATVs: Nothing to report
HabCar used and why, where? Yes, hauling water
CrewCar used and why, where? Yes, the team took it to the Goblin Valley this afternoon and it was used at the storage units this morning. It will need to be cleaned before crew use.
Luna used and why, where? No
General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. We will work on radios tomorrow.
Campus-wide inspection, if action taken, what and why?

Volunteer vegetation around campus and MOOP have been removed.
I don’t know if I mentioned yesterday that we outlined in rocks a parking spot for visitors and borrowed the URC orange cones to stop people from coming up to MDRS.
Summary of general operations: We began work on the flooded storage units today but removing anything that has not been used in several years. We announced around town that it was first come, first served on anything outside. We will go tomorrow afternoon and empty more.

The goal is to have everything on shelves, then power wash the mud out, and hopefully have time to move all remaining materials into a single storage unit with nothing on the floor.

Summary of Hab operations: I remembered the leak in the kitchen sink today because the ground is wet again. That will need to be repaired before the first crew arrives.

Summary of Outpost operations: Septic and intern trailer blackwater tanks will be pumped on Thursday.

We finished the construction of the new bed in the intern trailer. It has a double bed on the bottom and a twin on top. They have the best view on Mars.

Summary of GreenHab operations: 60% shade cloth installed. The soil was added to the herb garden, the two greens containers, one of the carrot bins, plus 12 two-gallon pots and 21 one-gallon pots. We made two water runs for the GreenHab today.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Too hot. I am trying to get the team to embrace working in the early morning. So far, they aren’t embracing!

Questions, concerns, supplies needed and requests: Nothing at this time.

Shannon M Rupert, PhD
Director
Mars Desert Research Station
P.O. Box 82
Hanksville, Utah 84734

Supplemental Operations Report 04-SEPT-2021

MDRS Supplemental Operations Report 04-SEPT-2021

This is the second of two operations reports filed tonight.

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party

Non-nominal systems: This is a reminder of what systems needed repair when we left: Leak in Hab toilet, leak in upper deck water tank intake, GreenHab door, leak in ScienceDome under window with damage

Action taken for non-nominal systems: These are all issues that need to be addressed before the field season begins

Generator check, note if oil and coolant added: New generator has not been activated for autostart. I will do a test run on the machine in the new two weeks.

ScienceDome Dual Split: Set to automatically keep temperature at 77 degrees F/24 degrees C for the summer, and worked perfectly. I turned it lower today so we could escape in there when it got hot in the afternoon.

Solar— Nominal, providing all power

Diesel Reading – 0 %.

Propane Reading, main tank – 80 %

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72% (loss of 3%) over summer

Generator propane tank: 80%

Water (Outpost tank) – 200 gallons

Campus wide inspection, if action taken, what and why? Contacted contractor to pump septic tank at Outpost. We will make arrangements for the main septic to be pumped later this fall. We created a rock outlined parking lot at the main sign as the first step in our tourist design.

Volunteer vegetation around campus is currently being removed.

Summary of Hab operations: There has been significant damage to the exterior shell, which we will be painting this week. In some places it has cracked and I am hopeful the elastomeric paint will seal it again. There is also some evidence of water around the window of the upper deck. We will try to attach drip flashing after painting, and repair interior damage as soon as we can. Thee staircase is very rusted as it was never properly finished and sealed. That is something that should be done between crews. At minimum rust needs to be removed.

Summary of Outpost operations: We attacked the intern trailer with a vengeance this morning! The sofa and dining booth were removed, as well as all the window valences. The designed and purchased material to build a full-size bunk-bed for the front living area and a wall to divide the bedroom from the bathroom to allow more privacy when a group of people are living there. The dining booth will be replaced with a table and chairs. The walls will be painted and new insulated curtains put on the windows. Additionally, we need to make repairs to the water system and repair the main furnace. Finally, the trailer is listing several degrees on one side, so we need to level it.

Summary of GreenHab operations: Nothing to report except the potting soil and some plants were purchased.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Some paint touch up should be done and some leaks need to be sealed. It was a very humid summer as we had some rusting of materials in the building.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: It’s been hotter and getting hotter so we are keeping the dome cool, drinking a lot of water and making sure not to work in the sun during the hottest time of the day.

Questions, concerns, supplies needed and requests: We have requested an additional person or two from the URC to come to Hanksville and help with the clean out of the storage units. As you may know, Hanksville was hit with heavy rain and there was major flooding in a concentrated area in town. That area included the storage facility where MDRS and the URC rent two units. The entire area was flooded and our units have between 2 and ½ inches of wet mud completely covering them. A lot of the Ikea materials for the ScienceDome were on the floor and had varying degrees of mud and water damage. A complete inventory will be needed to determine what we have to replace. In the meantime, we need to completely clear out the units and power wash them this week and get everything we can either out or cleaned up and put away. The remodel of the Science Dome has been put on hold for now. The large trailer will be brought to the storage place so we can get rid of stuff.

Operations Report 4-Sept-2021

Work Party Crew Operations Report 04-Sept-2021

This is one of two operations reports filed today

SOL: N/A

Name of person filing report: Shannon Rupert

Non-nominal systems: Intern trailer water, intern trailer heat.

Notes on non-nominal systems: Working on repairs

Generator: Nothing to report

Hours run: N/A

Solar— SOC % Last 24 hours :

Average 85.2

Minimum 61

Maximum 100

Station Propane Reading – 80 %

Ethanol Free Gasoline – not checked gallons

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 200 gallons

Static to Loft Pump used – yes

Water in GreenHab –0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: Yes

Hours: 216.4

Beginning charge: 96

Ending charge: 97

Currently charging: yes

Notes: When we first arrived, Percy was parked outside my trailer and plugged in but the malfunction light was flashing and it was at 18% when it should have been at 100%. We couldn’t determine why it was not charging so we moved it to the RAM and tried to charge it again. It is slowly charging and the light has stopped blinking. Later we discovered that the 110 plug in the outpost power tower was tripped, probably due to heavy rain. Rover appears to be nominal now.

Notes on rovers: The 4 remaining rovers have had their batteries replaced and were supposed to be delivered this morning, but there was a lack of communication between parties and they did not get delivered. My assumption is that delivery will be rescheduled for sometime after the holiday weekend.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). The three 350’s. The Honda is still non-operational. The 300 is no longer at MDRS.

Reason for use: Test for battery and road worthiness.

Oil Added? No

ATV Fuel Used: Not checked Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: The 350’s are all functional and in generally good operating order. We had to jump one, and tighten the battery cables on another but they all started. The tires were checked and air added as needed.

HabCar used and why, where? Not until this evening, when the new battery was installed. We will use it tomorrow to get water for the GreenHab,

CrewCar used and why, where? Yes, team took it to the Quarry and it was used to go check out the storage units in town.

Luna used and why, where? Yes, to Grand Junction for work party materials

General notes and comments: Grateful for all three vehicles today

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are as we left them in the station but without batteries, as the batteries were stored over summer in the ScienceDome and we have not put them back in the radios.

Summary of Hab operations: Mostly nominal. We have not turned on the hot water yet, but probably will tonight.

Summary of GreenHab operations: Potting soil was purchased We will add soil to the GreenHab. Fill the water tank and put up the shade cloth tomorrow. Seeds will be planted within the next couple of days.

Summary of ScienceDome operations: All nominal, no work done here.

Summary of RAM operations: This building really needs a good solid cleaning.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: We put out poison this summer to avoid rodent infestation, but most of it remained untouched. There was some evidence of pack rat visitors in the intern trailer and the Hab, but remaining buildings show no sign of rodents.

Questions, concerns and requests to Mission Support: Please add Marufa and Ashley to this eList. They are our Operations Interns for this season and they have been copied on this email.

We have requested additional personal from the URC team to assist with the cleaning and mud removal in the storage units.